OpenStudy (anonymous):

A piece of gold wire has a diameter of 0.175 cm. How much pricisely 1.00 x 10 to the fifth power of the wire weigh?

OpenStudy (anonymous):

@abb0t

OpenStudy (abb0t):

You want to find the weight of the wire with a diameter of 0.175 cm? Is that what you're asking?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are you given density? also the length of the wire?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its 19.3 g/cm^3 though

OpenStudy (anonymous):

ok, and i don't quite understand the second sentence. Could you clarify?

OpenStudy (anonymous):

how much does 1.0 x 10^5 cm of the gold wire weigh

OpenStudy (anonymous):

oh so that's the length... ok, so you have the diameter, you have the length so from there you can find the volume You also have the density, so once you found the volume, multiply it by the density and voila, mass!

OpenStudy (anonymous):

so the formula for volume is V=pi(d)(L)?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

that's the surface area of the wire

OpenStudy (anonymous):

volume = cross sectional area * length. Cross section area = pi * r^2 = pi * (d/2)^2

OpenStudy (anonymous):

also i forgot to add that youre supposed to imagine the wire to be a cylinder

OpenStudy (anonymous):

yea, i did. Math still doesn't change though

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

2405.281875 is what i got for volume

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

what next?

OpenStudy (anonymous):

volume x density = mass. So multiply them and that's your answer